# LV-UNet: A Lightweight and Vanilla Model for Medical Image Segmentation

### v3

@@ Abstract
Index Terms:
Medical image segmentation, lightweight, vanilla, MobileNetv3-Large, deep training strategy, re-parametrization
@@ IV-A Overall Architecture
Figure 1: The architecture of LV-UNet: the basic modules include pre-trained MobileNetv3-Large blocks(the initial convolution stage and the group \@slowromancapi@ to \@slowromancapiii@ (the first inverted residual block to ninth), fusible encoder blocks, fusible decoder blocks, skip-connections, and the output block.
@@ IV-B Pre-trained Modules in Encoder
LV-UNet uses MobileNetv3-Large from the initial convolutional stage to the ninth inverted residual block for the encoder block of LV-UNet. The initial convolutional stage receives the input image and performs basic feature extraction, and the following layers contain inverted residual blocks with depth-wise separable convolution layers, different convolution kernels and activation functions, progressively extracting more abstract features. We name inverted residual blocks in the MobileNetv3-Large model as $r1$ to $r15$. These inverted residual blocks are grouped as follows: $r1$ and $r2$ as the first group (group \@slowromancapi@) , $r3$, $r4$, and $r5$ as the second group (group \@slowromancapii@), and $r6$, $r7$, $r8$, and $r9$ as the third group (group \@slowromancapiii@). We do not use the entire feature extraction part of the MobileNetv3-Large model, to allow flexible design of subsequent expansion modules.
@@ IV-C2 Nonlinear Activation Layer
where $A_{s}\left(x_{h,w,c}\right)$ denotes the output of the activation function at position $(h,w,c)$ in the input tensor $h$ ranges from $1$ to $H$, $w$ ranges from $1$ to $W$, and $c$ ranges from $1$ to $C$. The function aggregates inputs from neighboring positions $(i+h,j+w,c)$ within a range defined by $-n$ to $n$. The weights $a_{i,j,c}$ control the contribution of each input, and $A(\cdot)$ represents the activation function applied to each input, typically a nonlinear function like ReLU. Furthermore, $b_{c}$ represents a bias term specific to each channel $c$. By utilizing this formulation, LV-UNet can capture and incorporate contextual information from neighboring positions, enabling the model to learn and represent complex patterns and nonlinear relationships within the input data.
@@ IV-C3 Re-parametrization and Deployment Mode
Then the two 1$\times$1 convolution layers in the fusible blocks can be merged as a single convolution layer:
@@ V-A4 CVC-ColonDB
The CVC-ColonDB dataset [16] includes 380 images, each with a resolution of 574$\times$500 pixels. These images were selected from a pool of 12,000 images obtained from 15 short colonoscopy videos, with only these 380 images being annotated. The test set is split with a ratio of 0.2. Then the rest of the dataset is split into the training set and the validation set with a ratio of 8:2.
@@ V-A5 Kvasir-SEG
The Kvasir-SEG dataset [17] comprises 1000 polyp images and their corresponding ground truth masks with resolutions ranging from 332$\times$487 to 1920$\times$1072 pixels. The test set is split with a ratio of 0.2. Then the rest of the dataset is split into the training set and the validation set with a ratio of 8:2.
@@ V-B Implementation Details
The experiments are all performed on PG500-216(V-100) with 32 GB memory. The total training epochs are 300, and the batch size in training, validation, and testing is 8. The resolution of the input images is resized to 256$\times$256. The optimizer used is ADAM [26]. The initial learning rate is 0.001, and a CosineAnnealingLR [27] scheduler is used. The minimum learning rate is 0.00001. Only rotation by 90 degrees clockwise for random times, random flipping, and normalization methods are used for data augmentation. The evaluation metric in the validation is $IOU$. The loss function used is a mixed loss that combined binary cross-entropy (BCE) loss and dice loss [28]:
@@ V-C Results
The experimental results of the comparison can be seen in Table \@slowromancapi@, using $IOU$ and $Dice$ as metrics, from which we can see that LV-UNet shows competitive performances compared to baselines. In addition, the deployment mode of LV-UNet can still achieve good results while reducing parameters a lot. In Table \@slowromancapii@, the number of parameters and FLOPs of different models are compared, which shows that the model sizes and computation complexities of LV-UNet and its deployment mode are suitable for edge device and point of care scenarios. Visualizations of some segmentation results can be seen in Fig.3.
@@
Combination \@slowromancapi@ 0.6308 0.7548 0.8531 0.9197 0.8063 0.8901
Combination \@slowromancapii@ 0.6489 0.7785 0.8514 0.9183 0.8069 0.8917
Combination \@slowromancapiii@ 0.6332 0.7625 0.8324 0.9080 0.7971 0.8856
@@
      Combination \@slowromancapi@       2.8M       0.34G
      Combination \@slowromancapii@       0.9M       0.22G
      Combination \@slowromancapiii@       0.8M       0.17G
@@ V-D Ablation Studies
    Combination \@slowromancapi@: From the initial convolutional stage to the inverted residual block $r14$ in MobileNetv3-Large and two fusible encoder blocks
@@
    Combination \@slowromancapii@: From the initial convolutional stage to the inverted residual block $r9$ in MobileNetv3-Large and three fusible encoder blocks
@@
    Combination \@slowromancapiii@: From the initial convolutional stage to the inverted residual block $r5$ in MobileNetv3-Large and four fusible encoder blocks
@@
The comparison results can be seen in Table \@slowromancapiii@, and the parameters and FLOPs can be seen in Table \@slowromancapiv@. The results show that the LV-UNet setting (combination \@slowromancapii@) can achieve better performance on BUSI and Kvasir-SEG datasets and achieve a very close result on CVC-ClinicDB dataset. Meanwhile, compared to combination \@slowromancapi@, combination \@slowromancapii@ reduces the model size and computational complexities.
The Number of Series-informed Activation Functions. In experiments in [21], the increase of the number of series-informed activation functions can lead to improvements in VanillaNet’s classification accuracy. However, in our experiments, no significant relationship was observed between the number of series-informed activation functions and the segmentation accuracy of LV-UNet. However, the results suggest that the use of an appropriate number of series-informed activation functions may potentially enhance segmentation accuracy. The comparison results can be seen in Table \@slowromancapv@.
@@
The Way of Skip-Connections. We evaluated the two types of skip-connections methods: adding and concatenation. The adding operation in skip-connections outperforms concatenation on all three datasets while reducing parameters and computational complexities. The comparison results can be seen in Table \@slowromancapvi@.
Deep Training Strategies. We also compared different nonlinearity decreasing methods: $a=1-cos({\pi e}/{2E})$ and the design in  [21] $a={e}/{E}$. The comparison results in Table \@slowromancapvii@ show that the performances are very close and that our method achieves better results in more datasets.

### v4

@@ IV-A Overall Architecture
Figure 1: The architecture of LV-UNet: the basic modules include pre-trained MobileNetv3-Large blocks(the initial convolution stage and the group I to III (the first inverted residual block to ninth), fusible encoder blocks, fusible decoder blocks, skip-connections, and the output block.
@@ IV-B Pre-trained Modules in Encoder
LV-UNet uses MobileNetv3-Large from the initial convolutional stage to the ninth inverted residual block for the encoder block of LV-UNet. The initial convolutional stage receives the input image and performs basic feature extraction, and the following layers contain inverted residual blocks with depth-wise separable convolution layers, different convolution kernels and activation functions, progressively extracting more abstract features. We name inverted residual blocks in the MobileNetv3-Large model as $r1$ to $r15$. These inverted residual blocks are grouped as follows: $r1$ and $r2$ as the first group (group I) , $r3$, $r4$, and $r5$ as the second group (group II), and $r6$, $r7$, $r8$, and $r9$ as the third group (group III). We do not use the entire feature extraction part of the MobileNetv3-Large model, to allow flexible design of subsequent expansion modules.
@@ IV-C2 Nonlinear Activation Layer
where $A_{s}\left(x_{h,w,c}\right)$ denotes the output of the activation function at position $(h,w,c)$ in the input tensor $h$ ranges from $1$ to $H$, $w$ ranges from $1$ to $W$, and $c$ ranges from $1$ to $C$. The function aggregates inputs from neighboring positions $(i+h,j+w,c)$ within a range defined by n-n to $n$. The weights $a_{i,j,c}$ control the contribution of each input, and $A(\cdot)$ represents the activation function applied to each input, typically a nonlinear function like ReLU. Furthermore, $b_{c}$ represents a bias term specific to each channel $c$. By utilizing this formulation, LV-UNet can capture and incorporate contextual information from neighboring positions, enabling the model to learn and represent complex patterns and nonlinear relationships within the input data.
@@ IV-C3 Re-parametrization and Deployment Mode
Then the two $\times1$ convolution layers in the fusible blocks can be merged as a single convolution layer:
@@ V-A4 CVC-ColonDB
[…]
@@ V-A5 Kvasir-SEG
[…]
@@ V-B Implementation Details
The experiments are all performed on PG500-216(V-100) with 32 GB memory. The total training epochs are 300, and the batch size in training, validation, and testing is 8. The resolution of the input images is resized to $\times256$. The optimizer used is ADAM [26]. The initial learning rate is 0.001, and a CosineAnnealingLR [27] scheduler is used. The minimum learning rate is 0.00001. Only rotation by 90 degrees clockwise for random times, random flipping, and normalization methods are used for data augmentation. The evaluation metric in the validation is $IOU$. The loss function used is a mixed loss that combined binary cross-entropy (BCE) loss and dice loss [28]:
@@ V-C Results
The experimental results of the comparison can be seen in Table I, using $IOU$ and $Dice$ as metrics, from which we can see that LV-UNet shows competitive performances compared to baselines. In addition, the deployment mode of LV-UNet can still achieve good results while reducing parameters a lot. In Table II, the number of parameters and FLOPs of different models are compared, which shows that the model sizes and computation complexities of LV-UNet and its deployment mode are suitable for edge device and point of care scenarios. Visualizations of some segmentation results can be seen in Fig.3.
@@
Combination I 0.6308 0.7548 0.8531 0.9197 0.8063 0.8901
Combination II 0.6489 0.7785 0.8514 0.9183 0.8069 0.8917
Combination III 0.6332 0.7625 0.8324 0.9080 0.7971 0.8856
@@
   Combination I    2.8M    0.34G
   Combination II    0.9M    0.22G
   Combination III    0.8M    0.17G
@@ V-D Ablation Studies
    Combination I: From the initial convolutional stage to the inverted residual block $r14$ in MobileNetv3-Large and two fusible encoder blocks
@@
    Combination II: From the initial convolutional stage to the inverted residual block $r9$ in MobileNetv3-Large and three fusible encoder blocks
@@
    Combination III: From the initial convolutional stage to the inverted residual block $r5$ in MobileNetv3-Large and four fusible encoder blocks
@@
The comparison results can be seen in Table III, and the parameters and FLOPs can be seen in Table IV. The results show that the LV-UNet setting (combination II) can achieve better performance on BUSI and Kvasir-SEG datasets and achieve a very close result on CVC-ClinicDB dataset. Meanwhile, compared to combination I, combination II reduces the model size and computational complexities.
The Number of Series-informed Activation Functions. In experiments in [21], the increase of the number of series-informed activation functions can lead to improvements in VanillaNet’s classification accuracy. However, in our experiments, no significant relationship was observed between the number of series-informed activation functions and the segmentation accuracy of LV-UNet. However, the results suggest that the use of an appropriate number of series-informed activation functions may potentially enhance segmentation accuracy. The comparison results can be seen in Table V.
@@
The Way of Skip-Connections. We evaluated the two types of skip-connections methods: adding and concatenation. The adding operation in skip-connections outperforms concatenation on all three datasets while reducing parameters and computational complexities. The comparison results can be seen in Table VI.
Deep Training Strategies. We also compared different nonlinearity decreasing methods: $a=1-cos({\pi e}/{2E})$ and the design in  [21] a=e/Ea={e}/{E}. The comparison results in Table VII show that the performances are very close and that our method achieves better results in more datasets.